# M4FC: a Multimodal, Multilingual, Multicultural, Multitask Real-World Fact-Checking Dataset

### v1

@@ Abstract
Existing real-world datasets for multimodal automated fact-checking have multiple limitations: they contain few instances, focus on only one or two languages and tasks, suffer from evidence leakage, or depend on external sets of news articles for sourcing true claims. To address these shortcomings, we introduce M4FC, a new real-world dataset comprising 4,982 images paired with 6,980 claims. The images, verified by professional fact-checkers from 22 organizations, represent diverse cultural and geographic contexts. Each claim is available in one or two out of ten languages. M4FC spans six multimodal fact-checking tasks: visual claim extraction, claimant intent prediction, fake detection, image contextualization, location verification, and verdict prediction. We provide baseline results for all tasks and analyze how combining intermediate tasks influence downstream verdict prediction performance. We make our dataset and code available.111github.com/UKPLab/M4FC
@@ 1 Introduction
Since 2020, approximately 80% of online misinformation verified by fact-checkers has been multimodal, combining text with images, videos, or audio (Dufour et al., 2024). Debunking multimodal misinformation is challenging and time-consuming for human fact-checkers (Silverman, 2013; Mossou and Higgins, 2021; Khan et al., 2023, 2024). It typically involves performing several sub-tasks and utilizing specialized tools like satellite imagery providers and sun angle calculators (Mossou and Higgins, 2021; Khan et al., 2023, 2024). Additionally, it requires expert knowledge of specific languages and cultural contexts (Silverman, 2013). This work specifically targets the two most common categories of image-based misinformation (Dufour et al., 2024): (1) out-of-context (OOC) misinformation, where authentic images are paired with false claims about their date, location, or the people and events depicted; and (2) false claims paired with manipulated or fake images.
Figure 1: An out-of-context instance from M4FC. The six AFC tasks are shown in blue and their output in gray.
Multimodal automated fact-checking (AFC) assists human fact-checkers by automating parts of their workflow (Akhtar et al., 2023). Several datasets have been proposed for multimodal AFC. Synthetic datasets, on the one hand, are large and scalable (Luo et al., 2021; Shao et al., 2024) but cannot fully mimic real-world misinformation. On the other hand, real-world datasets consist of claims verified by human fact-checkers, but these datasets tend to be much smaller (Zlatkova et al., 2019; Hu et al., 2023; Papadopoulos et al., 2024; Tonglet et al., 2024). Existing real-world datasets suffer from additional limitations. While multimodal misinformation is a global problem, these datasets cover only one or two languages and lack geographic diversity. Furthermore, the human fact-checking workflow involves several intermediate tasks before providing a verdict. However, the real-world datasets fail to account for most of these intermediate tasks.
@@ 2.1 Synthetic multimodal AFC datasets
Several synthetic datasets have been proposed for multimodal AFC (Sabir et al., 2018; Müller-Budack et al., 2020; Luo et al., 2021; Shao et al., 2024; Liu et al., 2025). While synthetic datasets are large and scalable, they often suffer from biases. Papadopoulos et al. (2025) demonstrated that simple models can exploit these biases to achieve near SOTA performance, raising concerns about the reliability of synthetic data for evaluating progress in multimodal AFC.
@@ 3.1 Six multimodal AFC tasks
Multimodal claims are not always expressed online in an explicit and verifiable format. Visual claim extraction (VCE) addresses this issue by generating an English-language claim from the screenshot of a social media post. This new task involves two challenges: (1) extracting visual cues and text from a wide range of languages, and (2) leveraging external knowledge. As shown in Figure 2, parts of the claim appear in the text, others on the building within the image, and further clarification requires external knowledge.
@@
Fake detection (FD) identifies images that are manipulated or fake. Manipulations include face swaps or background changes, while fake images encompass AI-generated content and human-made digital artwork. This classification task has two labels: authentic and manipulated/fake.
@@
Verdict prediction (VP) classifies a claim as true or false. The task is defined in four settings, based on two factors: whether claims are English-only or multilingual, and whether the verdicts are imbalanced or balanced. In the English-only setting, all claims are in English, while the multilingual setting covers 10 languages. The imbalanced setting captures the real-world distribution of fact-checking verdicts, in which the majority of claims are false. The balanced setting is constructed by replacing a portion of false claims with synthetic true claims linked to the same image.
@@ 3.2 Dataset construction
3.2.1 Data collection and labeling
We collected articles from 22 organizations, based in 17 different countries and all members of the International Fact-Checking Network (IFCN)222ifcncodeofprinciples.poynter.org/signatories, ensuring high geographic and linguistic diversity and adherence to professional standards. Table 2 reports the distribution of the images per fact-checking organization. Some organizations produce a higher volume of articles than others.
Following Tonglet et al. (2024), we use GPT4o (OpenAI, 2023) to extract claims in English and in the original language, task labels, and metadata from the article. GPT4o also generates synthetic true claims based on the context items of images labeled as false (out-of-context). For a random half of those instances, the original false claim is replaced by the synthetic true claim under the balanced setting. For the multilingual setting, we consider languages with at least 40 claims; claims in less-represented original languages are translated in English. All prompts are provided in Appendix A. Verdicts are mapped to the following categories: true, false, mixture, unverified. Finally, we apply a temporal split to the dataset to avoid temporal leakage (Glockner et al., 2022). Appendix B provides more details about the data collection and labeling.
@@ 3.2.2 Data validation
We adopt the validation procedure from Tonglet et al. (2024) to assess the quality of GPT4o annotations. For each organization, we randomly sample up to 20 articles. Each article is reviewed by two annotators who evaluate the correctness of the data fields extracted by GPT4o. Annotators are native speakers of the organization’s language and are recruited through Prolific.333prolific.com 79 annotators participated in the study, of which 33 were excluded for failing more than five out of 15 attention checks.
During post-processing of the Prolific results, we distinguish between incorrect and missing labels. A label is marked as missing when GPT4o outputs “not enough information”, despite relevant information being present in the article. A label is considered incorrect or missing only if both annotators agree on it. Under this criterion, the majority of labels are accurate, with average correctness per field ranging from 86.2% (for claimed date) to 99.31% (for verdict). Using a more lenient criterion, counting a label as incorrect if at least one annotator flags it, the correctness ranges from 49.6% (for claimed date) to 89.63% (for verdict), still indicating strong automated annotation quality.
We compute inter-annotator agreement (IAA) using Randolph’s $\kappa$ (Randolph, 2005), which ranges from 0.32 to 0.81 across all data fields, indicating moderate to strong agreement. If we exclude instances labeled by GPT4o as not “enough information”, the main source of disagreement, the IAA ranges from 0.45 to 0.85. Guidelines and detailed scores are provided in Appendices C.1 and C.2.
@@ 3.2.3 Dataset statistics
Table 3 provides the distribution of instances and the type of inputs per task. 1,148 images are unsuitable for visual claim extraction because they do not contain any embedded text, making it impossible to infer a claim. Location verification has fewer instances than other tasks, and may benefit from dedicated synthetic datasets in future work. Additional statistics are provided in Appendix D.
@@ 3.2.4 Evidence collection
For image contextualization and verdict prediction, we collect up to 20 web evidence using the Google RIS engine.444cloud.google.com/vision/docs/detecting-web For location verification, we collect maps from OpenStreetMap (OpenStreetMap contributors, 2017) and satellite images from ESRI World Imagery, 555arcgis.com/home/item.html accessed with staticmap.666github.com/komoot/staticmap
@@ 4.1 Evaluation metrics
For visual claim extraction and claimant intent prediction, we assess the generated texts using Rouge-L (Lin, 2004), METEOR (Banerjee and Lavie, 2005), and BERTScore (Zhang et al., 2020). For fake detection, location verification, and verdict prediction, we report the macro F1-score (F1). For image contextualization, following Tonglet et al. (2025), we adopt METEOR for source, motivation, things, and event, F1 for people, $\Delta$ for date, and $CO\Delta$ for location. $\Delta$ and $CO\Delta$ measure the inverse proportional distance from the ground-truth in years and in thousand kilometers, respectively.
@@ 4.2 Baselines
For fake detection, we evaluate three specialized detectors, Deep-Fake Detector (Prithiv, 2024), AI-Image Detector (Dafilab, 2025), and RealFake Detector (Mikedata, 2025), considering both their pre-trained versions and fine-tuned variants on the MF4C training set. For other tasks, we employ three widely adopted open-weight MLLMs, InternVL2.5-8B (Chen et al., 2024), Qwen2.5-VL-7B (Qwen-Team, 2025), and Llama-3.2-11B-Vision (MetaAI, 2024), alongside two proprietary models, GPT4o-mini (OpenAI, 2023) and Gemini-1.5-Flash (Gemini-Team, 2024), with task-specific prompts provided in Appendix E. In addition, for image contextualization and verdict prediction, we benchmark against the current state-of-the-art (SOTA) methods, COVE (Tonglet et al., 2025) and DEFAME (Braun et al., 2025), respectively. COVE combines two models: LlavaNext-7B (Liu et al., 2023, 2024) and Llama-3-8B (MetaAI, 2024). For DEFAME, we use GPT4o-mini as the backbone LLM. Appendix F provides implementation details.
@@ 4.3 Baselines results
We analyze results for all tasks on the test set. Additional result tables are provided in Appendix G.
@@
Figure 4: Visual claim extraction (VCE), claimant intent prediction (CIP), fake detection (FD), and verdict prediction (VP) results (%). All VP results are shown in the balanced setting.
@@
Figure 4(a) shows that most models achieve decent performance, with the exception of Llama-3.2-11B-Vision, while Gemini-1.5-Flash consistently outperforms all others. Rouge-L and METEOR scores suggest a strong alignment between generated and reference claims, whereas BERTScore indicates a more moderate level of semantic correspondence. We manually analyzed 50 randomly sampled predictions from Gemini-1.5-Flash. Among them, 38 claims are accurate, nine were partially correct but lacked external world knowledge, and three were incorrect due to the model’s inability to capture sarcasm in social media posts. Figure 5 presents two error cases: on the left, the model fails to connect the image with its broader context, although the claim publication date and the landscape hint at a connection with the Maui fires; on the right, the text in German translates to “There are over 100 years between the two photos. You can clearly see the ever-rising sea level, can’t you?”, yet the model overlooks the sarcasm and extracts a claim that directly contradicts the ground truth.
@@
Figure 4(b) shows that Gemini-1.5-Flash achieves the best performance, with a clear margin on BERTScore compared to other models. Llama-3.2-11B trails slightly behind. GPT4o-mini underperforms because it often refuses to execute the task, flagging claimant intent prediction as a potential policy violation. The overall scores are low, highlighting that real-world claimant intent prediction remains challenging for current MLLMs.
Fake detection
Figure 4(c) reports F1 scores for specialized fake detectors. While the pre-trained Deep-Fake detector achieves a relatively strong F1 of 70.31%, AI-Image detector is almost ineffective at only 6.41%, and RealFake detector remains modest at 60.17%. Although the M4FC train test is small and does not overlap in time with the test set, fine-tuning on it yields high performance gains across all models. This means that features can be learned from prior real-world cases of manipulated and fake images and generalized to newer cases.
@@
We evaluate MLLMs by comparing each claim image with candidate locations presented in three formats: map view, satellite view, or both. As a baseline, we also evaluate MLLM performance when candidate locations are provided in text form. As shown in Table 5, F1 is often lower with aerial views compared to the text baseline, in particular for commercial models. This suggests that many MLLMs can recognize the location but struggle to match the image to its corresponding aerial view. Model behaviors further diverge by modality: Llama-3.2-11B performs worse than random (49.18%) when given the candidate location as text, while Qwen2.5-VL-7B achieves its best scores with satellite imagery. Gemini-1.5-Flash consistently leads across all settings, and GPT4o-mini is the only model to benefit from combining map and satellite views. Despite these differences, overall F1 scores remain modest, underscoring the difficulty of this multi-image reasoning task. Error cases in Figure 6 illustrate common failure modes: the dam example shows the model’s difficulty in distinguishing fine-grained structural details, while the aqueduct example highlights limitations of using a fixed zoom level (15), as discussed further in Appendix H.
@@
Figures 4(d) compare models under both multilingual and English-only balanced settings. Proprietary models show strong linguistic robustness, maintaining comparable performance between multilingual and English claims, regardless of RIS evidence availability. By contrast, open-weight models display much larger disparities across languages. InternVL2.5-8B and Llama-3.2-11B, in particular, perform markedly worse on multilingual claims, with Llama-3.2-11B suffering the steepest drop. Interestingly, Qwen2.5-VL-7B shows the opposite trend, achieving slightly better results in multilingual settings. Across all systems, providing RIS evidence consistently improves results, and the effect is most pronounced for models struggling with multilingual inputs. These results highlight substantial variation in cross-lingual generalization: while leading proprietary models achieve near-linguistic parity, many open-weight alternatives remain challenged in multilingual contexts. Overall performance levels also underscore the difficulty of verdict prediction: no model exceeds 80% F1. The English-only performance of GPT4o-mini with DEFAME or with RIS evidence does not differ much. This suggests that RIS constitutes the most useful source of evidence, and other evidence search tools introduced by the DEFAME framework, like web search, image search, and geolocation, have a limited impact on performance.
@@
Table 6 presents a detailed evaluation of how intermediate task outputs affects verdict prediction. We consider both predicted and ground truth intermediate outputs, in the imbalanced, English-only setting. We use GPT4o Mini and Qwen2.5-VL-7B, the best performing MLLMs for verdict prediction (see Figures 4(d)). A case study is shown in Appendix I. Replacing ground truth claims with extracted ones substantially decreases F1, showing that inaccuracies in extracted claims harm downstream verdict prediction. Adding all or some of the other intermediate predicted outputs and the web evidence improves F1. Web evidence has the strongest impact. Surprisingly, using only predicted image contextualization, despite it relying on the same web evidence, performs worse than using the web evidence directly. This suggests that many errors are introduced during image contextualization, and that the web evidence is not sufficient to predict most context items during that intermediate step. When ground truth intermediate outputs are available, image contextualization yields the largest performance boost, improving F1 by 16.96 percentage points. In that case, adding web evidence provides little to no improvement. The highest F1 is achieved by combining the web evidence and all ground truth intermediate tasks, highlighting the complementarity of the intermediate tasks towards verdict prediction. These results show the promises of combining AFC tasks and open the door for future work on how to best implement these combinations, in terms of accuracy, interpretability, and computational efficiency.

### v2

@@ Abstract
Existing real-world datasets for multimodal fact-checking have multiple limitations: they contain few instances, focus on only one or two languages and tasks, suffer from evidence leakage, or rely on external sets of news articles for sourcing true claims. To address these shortcomings, we introduce M4FC, a new real-world dataset comprising 4,982 images paired with 6,980 claims. The images, verified by professional fact-checkers from 22 organizations, represent a diverse range of cultural and geographic contexts. Each claim is available in one or two out of ten languages. M4FC spans six multimodal fact-checking tasks: visual claim extraction, claimant intent prediction, fake image detection, image contextualization, location verification, and verdict prediction. We provide baseline results for all tasks and analyze how combining intermediate tasks influences verdict prediction performance. We make our dataset and code available.111github.com/UKPLab/M4FC
@@ 1 Introduction
Since 2020, approximately 80% of online misinformation verified by fact-checkers has been multimodal, combining text with images, videos, or audio (Dufour et al., 2024). Debunking multimodal misinformation is time-consuming for human fact-checkers (Silverman, 2013; Khan et al., 2024). It typically involves performing several sub-tasks and utilizing specialized tools, like satellite and street-view imagery (Mossou and Higgins, 2021; Khan et al., 2023, 2024). Additionally, it requires expert knowledge of specific languages and cultures (Silverman, 2013). This work targets the two most common categories of image-based misinformation (Dufour et al., 2024): (1) out-of-context (OOC) misinformation, where authentic images are misrepresented by false claims paired; and (2) misinformation using manipulated or fake images.
Figure 1: An out-of-context instance from M4FC. The six AFC tasks are shown in blue and their ground truth output in grey. Purple boxes indicate (partially) incorrect LLM answers, while green boxes indicate correct ones.
Multimodal automated fact-checking (AFC) assists human fact-checkers by automating parts of their workflow (Akhtar et al., 2023). Several datasets have been proposed for multimodal AFC. Synthetic datasets, on the one hand, are large and scalable (Luo et al., 2021; Shao et al., 2024) but cannot fully mimic real-world misinformation. On the other hand, real-world datasets consist of claims verified by human fact-checkers, but these datasets tend to be much smaller (Zlatkova et al., 2019; Hu et al., 2023; Papadopoulos et al., 2024; Tonglet et al., 2024). Existing real-world datasets suffer from additional limitations. While multimodal misinformation is a global problem, these datasets cover only one or two languages and lack geographic diversity. Furthermore, the human fact-checking workflow involves several intermediate tasks before providing a verdict. However, existing datasets fail to account for most of these tasks.
@@ 2.1 Synthetic multimodal AFC datasets
Several synthetic datasets have been proposed for multimodal AFC (Sabir et al., 2018; Müller-Budack et al., 2020; Luo et al., 2021; Shao et al., 2024; Liu et al., 2025). While synthetic datasets are large and scalable, they often suffer from biases. Papadopoulos et al. (2025) demonstrated that simple models can exploit these biases to achieve SOTA performance, raising concerns about the reliability of synthetic data for multimodal AFC.
@@ 3.1 Six multimodal AFC tasks
Claims are not always expressed online in an explicit and verifiable format. Visual claim extraction (VCE) generates an English-language claim from the screenshot of a social media post. This task involves two challenges: (1) extracting visual cues and text from a wide range of languages, and (2) leveraging external knowledge. As shown in Figure 2, parts of the claim appear in the text, others on the building, and further clarification requires external knowledge.
@@
Fake image detection (FD) identifies images that are manipulated or fake. Manipulations include face swaps or background changes, while fake images encompass AI-generated content and human-made digital artwork. This task has two labels: authentic and manipulated/fake.
@@
Verdict prediction (VP) classifies a claim as true or false. The task is defined in four settings, based on two factors: whether claims are English-only or multilingual, and whether the verdicts are imbalanced or balanced. The multilingual setting covers 10 languages. The imbalanced setting captures the real-world distribution of fact-checking verdicts, in which most claims are false. The balanced setting replaces a portion of false claims with synthetic true claims based on the image’s true context. The English-only imbalanced setting can be used for training, while other settings are only recommended for evaluation, due to the presence of linguistic biases, as discussed in Appendix A.
@@ 3.2 Dataset construction
3.2.1 Data collection
The data in M4FC is labeled by human fact-checking experts. We collected fact-checking articles from 22 organizations, based in 17 different countries, all of which are members of the International Fact-Checking Network (IFCN), ensuring adherence to professional standards. Table 2 reports the distribution of images per fact-checking organization. Appendix E provides the geographic distribution of the images. M4FC images span six continents and 159 countries, resulting in greater cultural diversity and reduced geographic bias compared to prior work. The top three continents represented are Asia, Africa, and Europe.
All task labels and metadata are provided in the human-written raw article text. Following Tonglet et al. (2024), we utilize GPT4o (OpenAI, 2023) to extract task labels and metadata from the article, if available. Afterwards, GPT4o generates additional synthetic true claims based on the context items of images labeled as false (out-of-context). For a random half of those instances, the original false claim is replaced by the synthetic true claim under the balanced verdict prediction setting. All prompts are provided in Appendix B.
Images suitable for location verification are identified semi-automatically. We use keywords to filter articles referencing maps, satellites, street names, or landmarks. Furthermore, the image needs to have a location context item that can be mapped to coordinates. We then manually verify that the image shows an outdoor scene. Finally, GPT4o generates a list of likely locations for the image, and the first two with valid coordinates serve as negative candidate locations.
Finally, we apply a temporal dataset split to avoid leakage (Glockner et al., 2022). Appendix C provides more details about data collection.
@@ 3.2.2 Data validation
For each organization, we randomly sample up to 20 articles. Following Tonglet et al. (2024), each article is reviewed by two annotators who evaluate the correctness of the data fields extracted by GPT4o. Annotators are recruited through the Prolific platform. 79 annotators participated in the study, of which 33 were excluded for failing more than five out of 15 attention checks. A label is considered incorrect if both annotators agree on it. The majority of labels are accurate, with scores ranging from 86.2% (for claimed date) to 99.31% (for verdict). We compute inter-annotator agreement (IAA) using Randolph’s $\kappa$ (Randolph, 2005), which ranges from 0.32 to 0.81 across all data fields, indicating moderate to strong agreement. If we exclude instances labeled by GPT4o as not “enough information”, the main source of disagreement, the IAA ranges from 0.45 to 0.85. Guidelines and detailed scores are provided in Appendices D.1 and D.2.
@@ 3.2.3 Dataset statistics
Table 3 provides the distribution of instances per task. 1,148 images are unsuitable for visual claim extraction because they do not contain any embedded text to infer the claim. Location verification has the fewest instances because it applies to a specific subset of images that show outdoor scenes. More statistics are provided in Appendix E. We release a high-quality subset with labels validated by all annotators, as explained in Appendix F.
@@ 3.2.4 Evidence collection
For image contextualization and verdict prediction, we collect up to 20 web evidence using the Google RIS engine.222cloud.google.com/vision/docs/detecting-web For location verification, we use the coordinates of the candidate locations to collect maps from OpenStreetMap (OpenStreetMap contributors, 2017) and satellite images from ESRI World Imagery,333arcgis.com/home/item.html accessed with staticmap.444github.com/komoot/staticmap
@@ 4.1 Evaluation metrics
For visual claim extraction and claimant intent prediction, we assess the generated texts using Rouge-L (Lin, 2004), METEOR (Banerjee and Lavie, 2005), and BERTScore (Zhang et al., 2020). We also adopt G-Eval Liu et al. (2023b), a normalized LLM-based metric aligned with human judgment, to evaluate consistency between predictions and ground truth. We use GPT4o as the LLM. Appendix G provides the G-Eval prompts. For fake image detection, location verification, and verdict prediction, we report the macro F1-score (F1). For image contextualization, following Tonglet et al. (2025), we adopt METEOR for source, motivation, things, and event, F1 for people, $\Delta$ for date, and $CO\Delta$ for location. $\Delta$ and $CO\Delta$ measure the inverse proportional distance from the ground-truth in years and in thousand kilometers, respectively. For verdict prediction with intermediate tasks, we report the accuracy (Acc), precision (Pre), recall (Rec) and F1.
@@ 4.2 Baselines
For fake image detection, we evaluate three specialized detectors, Deep-Fake Detector (Prithiv, 2024), AI-Image Detector (Dafilab, 2025), and RealFake Detector (Mikedata, 2025), considering both their pre-trained versions and fine-tuned variants on the MF4C training set. For other tasks, we employ three widely adopted open-weight MLLMs, InternVL2.5-8B (Chen et al., 2024), Qwen2.5-VL-7B (Qwen-Team, 2025), and Llama-3.2-11B-Vision (MetaAI, 2024), alongside two proprietary models, GPT4o-mini (OpenAI, 2023) and Gemini-1.5-Flash (Gemini-Team, 2024), with task-specific prompts provided in Appendix H. In addition, for image contextualization and verdict prediction, we benchmark against the current state-of-the-art (SOTA) methods, COVE (Tonglet et al., 2025) and DEFAME (Braun et al., 2025), respectively. COVE combines two models: LlavaNext-7B (Liu et al., 2023a, 2024) and Llama-3-8B (MetaAI, 2024). For DEFAME, we use GPT4o-mini as the LLM. Appendix I provides implementation details.
@@ 4.3 Baselines results
We analyze results for all tasks on the test set. Additional result tables are provided in Appendix J.
@@
Figure 4: Visual claim extraction (VCE), claimant intent prediction (CIP), fake image detection (FD), and verdict prediction (VP) results (%). All VP results are shown in the balanced setting.
@@
Figure 4(a) shows that most models achieve decent performance, with the exception of Llama-3.2-11B-Vision, while Gemini-1.5-Flash consistently outperforms all others. Rouge-L and METEOR scores show the token-level alignment between generated and reference claims, whereas BERTScore and G-Eval indicate a more moderate level of semantic correspondence. We manually analyzed 50 randomly sampled predictions from Gemini-1.5-Flash. Among them, 38 claims are accurate, nine were partially correct but lacked external world knowledge, and three were incorrect due to the model’s inability to capture sarcasm in social media posts. Figure 5 presents two error cases: on the left, the model fails to connect the image with its broader context, although the claim publication date and the landscape hint at a connection with the Maui fires; on the right, the model overlooks the sarcasm and extracts a claim that directly contradicts the ground truth.
@@
Figure 4(b) shows that Gemini-1.5-Flash achieves the best performance, with a clear margin on BERTScore and G-Eval compared to other models. Llama-3.2-11B trails slightly behind. GPT4o-mini underperforms because it often refuses to execute the task, flagging claimant intent prediction as a potential policy violation. The overall scores are low, highlighting that real-world claimant intent prediction remains challenging for current MLLMs.
Fake image detection
Figure 4(c) reports F1 scores for specialized fake detectors. While the pre-trained Deep-Fake detector achieves a relatively strong F1 of 70.31%, AI-Image detector is almost ineffective at only 6.41%. Although the M4FC train set is small and does not overlap in time with the test set, fine-tuning on it yields high performance gains across all models. This means that features can be learned from prior real-world cases and generalized to newer cases.
@@
We evaluate MLLMs by comparing each claim image with candidate locations presented in three formats: text, map, satellite, or both map and satellite. As shown in Table 5, F1 is often lower with aerial views compared to the text baseline, in particular for commercial models. This suggests that many MLLMs can recognize the location but struggle to match the image to its corresponding aerial view. Model behaviors further diverge by modality: Llama-3.2-11B performs worse than random (49.18%) when given the candidate location as text, while Qwen2.5-VL-7B achieves its best scores with satellite imagery. Gemini-1.5-Flash consistently leads across all settings, and GPT4o-mini is the only model to benefit from combining map and satellite views. Despite these differences, overall F1 scores remain modest, underscoring the difficulty of this multi-image reasoning task. Error cases in Figure 6 illustrate common failure modes: the dam example shows the model’s difficulty in distinguishing fine-grained structural details, while the aqueduct example highlights limitations of using a fixed zoom level (15), as discussed further in Appendix K.
@@
Figure 4(d) compares models under both multilingual and English-only balanced settings. Proprietary models show strong linguistic robustness, maintaining comparable performance between multilingual and English claims, regardless of RIS evidence availability. By contrast, open-weight models display much larger disparities across languages. InternVL2.5-8B and Llama-3.2-11B, in particular, perform markedly worse on multilingual claims, with Llama-3.2-11B suffering the steepest drop. Interestingly, Qwen2.5-VL-7B shows the opposite trend, achieving slightly better results in multilingual settings. Across all systems, providing RIS evidence consistently improves results, and the effect is most pronounced for models struggling with multilingual inputs. These results highlight substantial variation in cross-lingual generalization: while leading proprietary models achieve near-linguistic parity, many open-weight alternatives remain challenged in multilingual contexts. Overall performance levels also underscore the difficulty of verdict prediction: no model exceeds 80% F1. The English-only performance of GPT4o-mini with DEFAME or with RIS evidence does not differ much. This suggests that RIS constitutes the most useful source of evidence, and other evidence search tools introduced by the DEFAME framework, like web search, image search, and geolocation, have a limited impact on performance.
@@
Table 6 presents a detailed evaluation of how intermediate task outputs affects verdict prediction. We consider both predicted and ground truth intermediate outputs, in the imbalanced, English-only setting. We use GPT4o Mini and Qwen2.5-VL-7B, the best performing MLLMs for verdict prediction in the balanced setting (see Figures 4(d)). Figure 1 shows an example of pipeline predictions with GPT4o Mini. Replacing ground truth claims with extracted ones substantially decreases F1, showing that inaccuracies in extracted claims harm downstream verdict prediction. Adding all or some of the other intermediate predicted outputs and the web evidence improves F1. Web evidence has the strongest impact. Surprisingly, using only predicted image contextualization, despite it relying on the same web evidence, performs worse than using the web evidence directly. This suggests that the web evidence is not sufficient to predict accurately most context items during the intermediate image contextualization step. When ground truth intermediate outputs are available, image contextualization yields the largest performance boost, improving F1 by 16.96 percentage points. In that case, adding web evidence provides little to no improvement. The highest F1 is achieved by combining the web evidence and all ground truth intermediate tasks, highlighting the complementarity of the intermediate tasks towards verdict prediction. These results show the promises of combining AFC tasks and open the door for future work on how to best implement these combinations, in terms of accuracy, interpretability, and computational efficiency.
We compare these results with the end to end system DEFAME. DEFAME’s F1 remains lower than the pipeline with ground truth intermediate task outputs. However, it largely outperforms the pipeline with predicted intermediate outputs. This shows that, while intermediate tasks are useful and necessary to produce a fact checking article, SOTA end to end systems like DEFAME still have a clear advantage for verdict prediction.